OpenStudy (meilendurcer):

–6w < 84 A. w > 14 B. w < –14 C. w > –14 D. w < 14

OpenStudy (anonymous):

-14

OpenStudy (meilendurcer):

@mark_o. can you help?

OpenStudy (anonymous):

c

OpenStudy (meilendurcer):

thanks

OpenStudy (anonymous):

no problem and can i have a medal

OpenStudy (meilendurcer):

yup :)

OpenStudy (meilendurcer):

@mjneedshelp thanks, can you help me with others?

OpenStudy (anonymous):

sure

OpenStudy (meilendurcer):

5m > 45 A. m > 9 B. m < 9 C. m > –9 D. m < –9

OpenStudy (anonymous):

a

OpenStudy (meilendurcer):

\[\frac{ n }{ -3 } \ge 4\] A. n ≤ 12 B. n ≥ –12 C. n ≤ –12 D. n ≤ 12

OpenStudy (anonymous):

c

OpenStudy (meilendurcer):

\[\frac{ k }{ 6 } > 4\] A. k < 24 B. k < –24 C. k > –24 D. k > –24